OpenStudy (anonymous):

Translate the sentence into an equation using n to represent the unknown number. Then solve the equation for n. the quotient of a number divided by 11 is 0.5

OpenStudy (anonymous):

Translate the following sentence into an equation using n to represent the unknown number. Then solve the equation for n. six more than twice a number is 16 a. 6(2 + n) = 16 n = 5 b. 6(2 + n) = 16 n = 4072-01-01-03-00_files/i0310000.jpg c. 6 + 2n = 16 n = 5 d. 6 + 2n = 16 n = 10